OpenStudy (anonymous):

A spaceship is orbiting the Earth around the circular orbit of radius R=2.5Ro, where Ro is the Earth’s radius. What should be the additional velocity V to be supplied to the spaceship in the direction away from the center of the Earth (along the radius of the orbit) so the ship may completely escape the Earth gravity?

OpenStudy (anonymous):

here is the picture

OpenStudy (anonymous):

I think you would need the mass of the spaceship to determine this, unless i am missing something.

OpenStudy (anonymous):

An energy perspective is most useful here. You need sufficient kinetic energy directed radially outward to raise the potential energy of the spacecraft from what it is at R to what it would be at infinity. \[dU = 0 - \left( - \frac{G M m}{R} \right) = \frac{G M m}{R}\] Set that equal to the additional kinetic energy: \[dE_k = \frac{1}{2} m v^2 = \frac{G M m}{R}\] Now solve for v. G is Newton's constant and M is the mass of the Earth. Fortunately the mass of the spacecraft drops out, so you don't need to know it.